OpenStudy (anonymous):

MAJOR HELP PLEASE!! Verify that the equation is an identity. Show all your work. (sin x + cos x)^2 - 1 = sin 2x

OpenStudy (nubeer):

(a+b)^2 =a^2+2ab+b^2 so (sin x + cos x)^2 - 1 sin^2x+2sinxcosx+cos^2x -1 2sinxcosx +sin^2x+cos^2x -1 i think u can do from here.

OpenStudy (tamtoan):

use formula \[(a+b)^{2} = a^{2} + 2ab + b^{2}\], expand that out and you will see what you need to

OpenStudy (anonymous):

@nubeer i'm sorry i am a little lost, could you please explain further?

OpenStudy (nubeer):

2sinxcosx +sin^2x+cos^2x -1 have you understood how i have got this part?

OpenStudy (anonymous):

yes with the formula right?

OpenStudy (nubeer):

2sinxcosx +sin^2x+cos^2x -1 what is sin^2x+cos^2x = ?

OpenStudy (anonymous):

1?

OpenStudy (nubeer):

yes.. now put this 1 in the equation 2sinxcosx +sin^2x+cos^2x -1 2sinxcosx +1-1.. got it?

OpenStudy (nubeer):

@kenneyfamily so?

OpenStudy (anonymous):

2sinxcosx

OpenStudy (anonymous):

@nubeer right?

OpenStudy (nubeer):

yes and 2sinxcosx=sin2x (according to the identity)

OpenStudy (anonymous):

you're awesome..thanks!!!